x²-6x-7= What are the factors of x²-5x, x-4 and x+6
Answer by stanbon(75887)

x²-6x-7 = (x-7)(x+1)
------------------------------
What are the factors of
x²-5x = x(x-5)
x-4 prime
x+6 prime
================
Cheers,
Stan H.
---------------